Max,Mon and Mitch are cousins and the sum of their ages is 64. Find their ages if Max is 3 times as old as Mitch and Mon's age is 4 more than twice the age of Mitch.
Answer by addingup(3677)

Ma+Mo+Mi = 64
Ma= Mi x 3
Mo= 2Mi+4 Substituting in the 1st formula:
(Mix3)+2Mi+4+Mi= 64 Subtract 4 from both sides and add the Mi on left:
Mix3+3Mi= 60 To make it easier, we break down the multiplication:
Mi+Mi+Mi+3Mi= 60
6Mi= 60
Mi= 10
Answer:
Mi is 10
Ma is 10x3= 30
Mo is 2(10)+4= 24
Proof:
10+30+24= 64 We've got the correct answer.